OpenStudy (anonymous):

Evaluate the function for the given value of X f(x)=|x-2| when x=5

OpenStudy (anonymous):

just substitute the value of x to f(x)

OpenStudy (anonymous):

like f(5)=|5-2| ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay i got 3

OpenStudy (anonymous):

that should be the answer.

OpenStudy (anonymous):

wait is that right?

OpenStudy (anonymous):

yes. why? is it showing you otherwise?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how many answers is it asking? cuz there is also the (-) side of it which would give you -3

OpenStudy (anonymous):

oh nevermind thats right thanks :)

OpenStudy (anonymous):

(x -2) = y or -(x-2) = y (5 - 2) = y or -(5-2) = y 3 = y or -3 = y

OpenStudy (anonymous):

ok cool